A gas that has a volume of 28 liters, a temperature of 45 °C, and an unknown pressure, has its volume increased to 34 liters and its temperature decreased to 35 °C. If I measure the pressure after the change to be 2.0 atm, what was the original pressure of the gas?
1
Expert's answer
2012-12-24T10:42:04-0500
For answering this question,you need to usenext equation

P1V1/T1 = P2V2/T2
P1 = P2V2T1/T2V1
P1 = 2.0 * 34 * 45 / 35* 28 = 3.12 atm